OpenStudy (anonymous):

Pump 1 takes 7minutes to fill my 20 gallon gas tank. Pump 2 takes 13 minutes to fill my tank. How long would it take if I could use both tanks at the same time?

OpenStudy (dumbcow):

\[\frac{20}{7}x + \frac{20}{13}x = 20\] add fractions 7*13 = 91 \[\frac{(20)(13)x + (20)(7)x}{91} = 20\] multiply 91 on both sides add like terms \[400x = 1820\] \[x = 4.55\]

OpenStudy (anonymous):

ok i understand the answer but not understanding the directions plz make them clearer

OpenStudy (anonymous):

is answer coorect?

OpenStudy (anonymous):

dont know

OpenStudy (dumbcow):

for pump 1 we need to find the rate of gall/min ->20/7 pump 2 rate is 20/13 gal/min if we multiply this rate by time we get a num of gallons this time is represented by x, the amount of time to fill up using both pumps total num of gallons needed is 20 so it equals 20

OpenStudy (anonymous):

ok gotcha

OpenStudy (anonymous):

i just have to explain in detail how i got my answer and show my work just alittle confused on your work